# https://www.recommender-systems.com Green Recommender Systems: Optimizing Dataset Size for Energy-Efficient Algorithm Performance

### v1

@@ Abstract
Reference: Arabzadeh, Ardalan; Vente, Tobias; Beel, Joeran. 2024. Green Recommender Systems: Optimizing Dataset Size for Energy-Efficient Algorithm Performance. In: International Workshop on Recommender Systems for Sustainability and Social Good (RecSoGood) at the 18th ACM Conference on Recommender Systems (ACM RecSys).
@@ 1 Introduction
Given the environmental and computational challenges associated with large datasets, it’s important to question whether using the entire datasets is always necessary. For instance, datasets like MovieLens 10M are frequently employed in training recommender systems. But is it necessary to use 10 million instances, especially with simple baseline algorithms? Or would downsampling the dataset to e.g. 10% suffices, which, in turn, might save 90% of energy? Similarly, dataset size might be an important factor when choosing datasets for recommender-system experiments [10].
@@ 2 Related Work
The field of green recommender systems has only started evolving recently [30, 35]. We also recently proposed "e-fold cross-validation", an energy-efficient alternative to k-fold cross-validation [11, 23, 5]. Wegmeth et al. introduced EMERS, a tool to measure the electricity consumption of recommender system experiments [36]. Also, judging on the "accepted papers" list of the RecSoGood workshop, more related work is to be published very soon [26, 32].
@@
These studies underscore and evaluate the potential benefits of downsampling and its impact on model performance, but are not directly applicable to traditional recommender system algorithms, where such effects remain underexplored. Especially when thinking of automated recommender systems (AutoRecSys [19, 3, 34]), where large spaces of configurations must be searched, energy efficiency is fundamentally important.
@@ 4.0.3 Conclusion
Therefore, we answer our research question by affirming that it can be possible to identify an optimal trade-off between maintaining algorithmic performance and reducing dataset size. Specifically, our analysis shows that strategic downsampling may improve energy efficiency while maintaining performance comparable to the original dataset size, thereby supporting the optimization of AI systems and recommenders. However, more research is necessary to find out when exactly downsampling is a sensible approach, as sometimes, performance varies notably. We hope that in the long term, downsampling datasets becomes an accepted best-practice [8, 7], for the recommender-system community that helps contributing to green and sustainable recommender systems.

### v2

@@ 1 Introduction
Given the environmental and computational challenges associated with large datasets, it’s important to question whether using the entire datasets is always necessary. For instance, datasets like MovieLens 10M are frequently employed in training recommender systems. But is it necessary to use 10 million instances, especially with simple baseline algorithms? Or would downsampling the dataset to e.g. 10% suffices, which, in turn, might save 90% of energy?
@@ 2 Related Work
The field of green recommender systems has only started evolving recently [26, 30]. We also recently proposed "e-fold cross-validation", an energy-efficient alternative to k-fold cross-validation [19, 4]. Wegmeth et al. introduced EMERS, a tool to measure the electricity consumption of recommender system experiments [31]. Also, judging based on the "accepted papers" list of the RecSoGood workshop, more related work is to be published very soon [22, 28].
@@
These studies underscore and evaluate the potential benefits of downsampling and its impact on model performance, but are not directly applicable to traditional recommender system algorithms, where such effects remain underexplored.
@@ 4.0.3 Conclusion
Therefore, we answer our research question by affirming that it can be possible to identify an optimal trade-off between maintaining algorithmic performance and reducing dataset size. Specifically, our analysis shows that strategic downsampling may improve energy efficiency while maintaining performance comparable to the original dataset size, thereby supporting the optimization of AI systems and recommenders. However, more research is necessary to find out when exactly downsampling is a sensible approach, as sometimes, performance varies notably. We hope that in the long term, downsampling datasets becomes an accepted best-practice [7, 6], for the recommender-system community that helps to contribute to green and sustainable recommender systems.